OpenStudy (anonymous):

Help please! Could someone walk me through this one : ) (3x =1)^3/2 = 64

OpenStudy (anonymous):

Is this 3x-1 or 3x+1 on the inside of the parentheses? (or neither?)

OpenStudy (anonymous):

oops.. +

OpenStudy (anonymous):

(3x+1)^3/2=64 3x+1=64^2/3=16 3x=15 x=5

OpenStudy (anonymous):

I can see how you did it in my head but on paper what do I do with the 3/2?

OpenStudy (anonymous):

You raise both sides of the equation to 2/3 to cancel it out on the left.

OpenStudy (anonymous):

ok : )

OpenStudy (anonymous):

p.s. Thanks so much!!